A drug is eliminated from the body through urine. Suppose that for an initial dose of 50 milligrams, the amount A(t) in the body t hours later is given by A(t) = 50(0.9)t.
(a) Estimate the amount of the drug in the body 6 hours after the initial dose.
(b) What percentage of the drug still in the body is eliminated each hour?

Answer by Alan3354(69443)

A drug is eliminated from the body through urine. Suppose that for an initial dose of 50 milligrams, the amount A(t) in the body t hours later is given by A(t) = 50(0.9)t.
(a) Estimate the amount of the drug in the body 6 hours after the initial dose.
A(t) = 50*0.9t = 45t
A(6) = 45*6 = 270 mg ***** It's increasing ******
-------
(b) What percentage of the drug still in the body is eliminated each hour?
None, it increases.